Bramble, Inc. manufactures two products: missile range instruments and space pressure gauges. During April, 50 range instruments and 200 pressure gauges were produced, and overhead costs of $85,060 were estimated. An analysis of estimated overhead costs reveals the following activities.

Activities

Cost Drivers

Total Cost

1. Materials handling Number of requisitions

$36,000

2. Machine setups Number of setups

26,265

3. Quality inspections Number of inspections

22,795

$85,060


The cost driver volume for each product was as follows.

Cost Drivers

Instruments

Gauges

Total

Number of requisitions 385 615 1,000
Number of setups 205 310 515
Number of inspections 235 250 485
Determine the overhead rate for each activity.

Overhead Rate

Materials handling $

?   

per requisition
Machine setups $

?

per setup
Quality inspections $

?

per inspection

LINK TO TEXT

Assign the manufacturing overhead costs for April to the two products using activity-based costing. (Round per unit answers to 2 decimal places, e.g. 12.25.)

Instruments

Gauges

Total cost assigned $

?

$

?

Overhead cost per Unit $

?

$

?
Answer #1
1
The overhead rate for each activity.
Overhead Rate
//img.homeworklib.com/questions/45f6f720-78f1-11ea-bb61-57ac9146885e.gif?x-oss-process=image/resize,w_560
Materials handling
$36.00 per requisition
($36,000 / 1000)
Machine setups $51.00 per setup
//img.homeworklib.com/questions/45f6f720-78f1-11ea-bb61-57ac9146885e.gif?x-oss-process=image/resize,w_560
($26,265 / 515)
//img.homeworklib.com/questions/45f6f720-78f1-11ea-bb61-57ac9146885e.gif?x-oss-process=image/resize,w_560
Quality inspections
$47.00 per inspection
($22,795 / 485)
2 Instruments Gauges
Materials handling $13,860.00 $22,140.00
(385 x $36), (615 x $36)
Machine setups $10,455.00 $15,810.00
(205 x $51), (310 x $51)
Quality inspections $11,045.00 $11,750.00
(235 x $47), (250 x $47)
Total cost assigned $35,360.00 $49,700.00
50 200
Overhead cost per Unit $707.20 $248.50